# Block-to-Scene Pre-training for Point Cloud Hybrid-Domain Masked Autoencoders

### v1

@@ 4.3 Ablation Study
4.3.1 The Impact of Stop Gradients.
In our implementation, stopping gradients is crucial. This approach allows different encoders within the model to learn scene-level and object-level representations independently during block-to-scene pretraining. By ensuring that each encoder focuses solely on its specific task, we enhance the model’s generalization capability. Consequently, maintaining the accuracy and independence of each encoder’s learning objective during training is essential.
Since the position regression objective and the masked reconstruction objective are two distinct tasks in our pretraining process, failing to decouple the learning processes of the different encoders could lead to catastrophic forgetting. For instance, without gradient stopping, gradients from the object-level reconstruction tasks could backpropagate into the scene’s encoder, interfering with its ability to learn scene-level knowledge. By applying gradient stopping, we effectively prevent this interference, ensuring that each encoder remains focused on its specific task and thereby avoiding catastrophic forgetting.
To further validate this issue, we conducted experiments where we removed the Stop Gradients operation from our pipeline and retrained the PointHDMAE model. We then assessed its performance on both scene-level point cloud detection tasks and object-level classification tasks. This comparison allowed us to observe the impact of gradient stopping on the model’s ability to effectively learn and generalize across different tasks. As shown in table 4.3.1, after removing the Stop Gradients operation, there was a 8.6% decrease in AP50 for the scene-level detection tasks, clearly indicating a deterioration in the representation capabilities of the scene encoder. Similarly, there was a noticeable decline in performance across classification tasks. These results collectively demonstrate that decoupling the representation learning of different encoders is essential to avoid catastrophic forgetting.
@@
Object-Level Classification OBJ-BG OBJ-ONLY w/o Stop Gradients 93.29 92.43 w/ Stop Gradients 95.18 93.12 Scene-Level Detection $AP_{25}$ $AP_{50}$ w/o Stop Gradients 63.1 41.3 w/ Stop Gradients 66.8 49.9
Table 8: The impact of the number of point blocks.
Number of blocks FLOPs(G) PB-T50-RS 1 22 88.78 8 59 89.10 16 102 89.52 32 187 90.01 64 359 90.12
4.3.2 The Impact of the Number of Point Blocks.
The number of randomly selected point blocks during the pretraining phase has a significant impact on the representation learning of the object encoder. Each point block serves as a sample to train the object encoder. The more point blocks selected from a scene, the richer the knowledge the object encoder can learn. However, this also leads to a significant increase in computational complexity. Therefore, we need to select an appropriate number to achieve a balance between efficiency and performance.
We selected 1, 8, 16, 32, and 64 blocks respectively and pre-trained these models from scratch. We reported the computational floating-point operations (FLOPs) required for pre-training and the performance of the resulting object models. The trained models were fine-tuned on the PB-RS-T50 variant of ScanObjectNN, using its split to evaluate model performance. As illustrated in the table 4.3.1, our model’s computational complexity significantly increases with the number of blocks, while the performance of the model gradually improves and eventually levels off. To achieve a balance between performance and efficiency, we chose 32 blocks for our experiments.
4.4 Conclusion
In this paper, we first propose a point cloud hybrid-domain masked autoencoders model to address the generalization limitations of existing domain-specific models. Our hybrid model selectively activates the object encoder to handle object domain point clouds specifically and leverages these object encoders to assist the scene encoder in processing scene domain point clouds. Then, we propose a block-to-scene pre-training strategy to train our PointHDMAE model. This strategy involves joint training through random object mask reconstruction and position regression within the scene, enabling our domain-specific encoder models to learn general representations relevant to their respective domains. Finally, we demonstrated the generalization and superiority of our model through extensive experiments across various datasets and tasks from different domains.

### v3

@@ 4.4 Ablation Study
4.4.1 The Impact of Different Pretraining Strategies
Scene B. R. Object M. R. Coordinate S. T. Joint T. Classification Detection
88.27 45.8
88.32 41.2
88.86 41.2
88.32 43.8
88.86 43.8
87.21 47.4
89.14 48.8
Table 5: The Impact of Different Pretraining Strategies.
We first analyze the impact of each key component in the block-to-scene pretraining strategy. Our block-to-scene pretraining strategy consists of four key parts: scene-level block position regression (Scene B. P.), object-level block mask reconstruction (Object M. R.), coordinate space transformation (Coordinate S. T.) of object blocks (from scene space to object space), and joint pertaining (Joint T.) of objects and scenes (connecting objects and scenes via the Projection module in Figure 3). A detailed analysis of these components is provided in Table 5.
When using only Scene B. P. or Object M. R., the pretraining strategy enhances only the subsequent scene tasks or object tasks, failing to achieve simultaneous improvements. Even when both strategies are employed without Joint T., while simultaneous enhancement is possible, the inability of the two pipelines to interact prevents the complementary knowledge from being effectively learned, thereby limiting the representation capability. Only by jointly employing both regression and reconstruction can the most comprehensive representation be achieved.
Secondly, Coordinate S. T. is crucial for learning object representations. This is because the coordinates of blocks directly cropped from the scene belong to the scene space and are often not normalized within the range of [-1, 1]. In contrast, most object-level point cloud analysis assumes coordinates within the range of [-1, 1]. Thus, transforming the coordinates is essential. As shown in the table, omitting this transformation results in performance degradation on object tasks. This is because the scene-space coordinates disrupt the learning of object representations, highlighting the importance of this normalization step for effective object-level analysis.
4.4.2 The Impact of Stop Gradients
Methods Object Classification Scene Detection
ModelNet40 ScanObjectNN AP25 AP50
w/o S. G. 93.4 87.45 62.7 41.5
w/ S. G. 94.0 89.14 66.2 48.8
@@ 4.4.2 The Impact of Stop Gradients
In our implementation, gradient stopping plays a crucial role. This mechanism allows different encoders within the model to jointly learn domain-specific representations during block-to-scene pretraining without interfering with each other. Thus, ensuring the accuracy and independence of each encoder’s learning objectives during training is of paramount importance.
Since the position regression objective and the masked reconstruction objective are two distinct tasks in our pretraining process, failing to decouple the learning processes of the different encoders could lead to catastrophic forgetting. For instance, without gradient stopping, gradients from the object-level reconstruction tasks could backpropagate into the scene’s encoder, interfering with its ability to learn scene-level knowledge. By applying gradient stopping, we effectively prevent this interference, ensuring that each encoder remains focused on its specific task and thereby avoiding catastrophic forgetting. Table 6 reports the impact of applying gradient stopping. When the gradient stopping mechanism is removed, the AP50 on scene-level detection tasks drops by 7.3%, clearly indicating that the learning process of the scene encoder is disrupted. Similarly, significant performance degradation is observed in classification tasks. These results collectively demonstrate that decoupling the representation learning of different encoders is essential for preventing catastrophic forgetting.
4.4.3 The Impact of Object Expert Initialization
Methods Object Classification Scene Detection
ModelNet40 ScanObjectNN AP25 AP50
Scratch 93.2 88.38 65.7 48.5
Pre-trained Point-MAE 94.0 89.14 66.2 48.8
Table 7: The impact of object expert initialization.
We further analyze the impact of initializing the object expert model. In our implementation, we initialize the model using an object model pre-trained on ShapeNet, after which no object data is used. This initialization ensures that we can learn representations from the object domain prior. We further examine the effect of not using object priors (Scratch), where no object domain data is used during training. The results, as shown in Table 7, indicate that the absence of object priors has a minimal impact on scene representation learning, resulting in only slight performance degradation. However, it significantly affects the learning of object point cloud representations. For the real point cloud ScanObjectNN, even when trained from scratch, the model eventually surpasses the one trained only on ShapeNet55 (88.27). However, for the synthetic point cloud ModelNet40, there is a larger gap compared to the model trained only on ShapeNet55 (93.8). This is because the point blocks cut from the scene lack specific semantics, making their distribution more similar to real point clouds but significantly different from the distribution of synthetic point clouds, leading to poorer performance on synthetic point clouds.
5 Conclusion
In this paper, we propose a block-to-scene pretraining strategy to train a mixture-of-domain-experts model that seamlessly integrates scene domain and object domain experts. This strategy enables the simultaneous acquisition of object-level and scene-level representations, enabling the learning of comprehensive 3D representations. Extensive experiments conducted across various datasets and tasks validate the superiority and transferability of our proposed approach. Our findings highlight the potential of leveraging cross-domain knowledge to achieve robust and versatile 3D representation learning, paving the way for future study.
Acknowledgements
This work is supported in part by the National Natural Science Foundation of China, under Grant (62302309,62171248), Shenzhen Science and Technology Program (JCYJ20220818101014030,JCYJ20220818101012025), and the PCNL KEY project (PCL2023AS6-1).